OpenStudy (anonymous):

What values of x and y satisfy x + 2y=6 and 2x + 3y=4

OpenStudy (anonymous):

x + 2y=6 and 2x + 3y=4 x + 2y=6 eqn 1 multiply by -3 2x + 3y=4 eqn 2 multiply by 2

OpenStudy (texaschic101):

or, if it is easier for you, this can be solved by substitution. x + 2y = 6 x = -2y + 6 now just sub -2y + 6 in for x in the other equation and solve for y. Once y is found, sub that back into either of the original equations to find x.